OpenStudy (toxicsugar22):

solve 11m^2-14m=0 and show me the steps to it

OpenStudy (toxicsugar22):

never mind i got it

OpenStudy (anonymous):

Okay. So you have m in both terms. So you can factor out an m, which leaves you with the following equation: m(11m-14)=0 One of the solutions is now trivial, being m = 0. The other solution requires that you take the inside of the parentheses. 11m-14 = 0. This becomes 11m = 14 Dividing both sides by 11 gives you m = 14/11 So your solutions are: x=0, x=14/11